### v2

# MACTAS: Self-Attention-Based Module for Inter-Agent Communication in Multi-Agent Reinforcement Learning

Abstract.
Communication is essential for the collective execution of complex tasks by human agents, motivating interest in communication mechanisms for multi-agent reinforcement learning (MARL). However, existing communication protocols in MARL are often complex and non-differentiable. In this work, we introduce a self-attention-based communication module that exchanges information between the agents in MARL. Our proposed approach is fully differentiable, allowing agents to learn to generate messages in a reward-driven manner. The module can be seamlessly integrated with any action-value function decomposition method and can be viewed as an extension of such decompositions. Notably, it includes a fixed number of trainable parameters, independent of the number of agents. Experimental results on the SMAC and SMACv2 benchmarks demonstrate the effectiveness of our approach, which achieves state-of-the-art performance on a number of maps.
Key words and phrases:
Multi-Agent Reinforcement Learning, Cooperation in Multi-Agent Reinforcement Learning, Communication in Multi-Agent Reinforcement Learning
1. Introduction
Multi-agent reinforcement learning (MARL) Ning and Xie (2024); Du and Ding (2021); Nguyen et al. (2020); Oroojlooy and Hajinezhad (2023); Wong et al. (2023) promises efficient control in large-scale systems through the collective learning of many agents in dynamic environments. Multiple agents interact within a shared environment to learn optimal policies. MARL is defined as the process in which agents optimize their behaviors based on the history of their interactions with the environment, i.e., on their experience. Research in this area aims to design methods that enable agents to learn quickly, i.e., from minimal experience. The challenges and complexities in using MARL for controlling multi-agent systems include coordination among agents, non-stationary environments due to evolving agent policies, and scalability issues. Advanced MARL approaches often involve sophisticated techniques, including decentralized training, communication protocols among agents, and hierarchical methods to manage multi-agent interactions as effectively as possible. Theoretical advancements and practical applications of MARL demonstrate its potential in diverse fields ranging from robotics Orr and Dutta (2023) and autonomous driving Yadav et al. (2023); Schmidt et al. (2022); Althamary et al. (2019) to telecommunications Li et al. (2022b); Feriani and Hossain (2021).
Figure 1. 3s5z_vs_3s6z scenario in the StarCraft Multi-Agent Challenge (SMAC).
The most prominent research topics in MARL are the action-value function decompositions Wang et al. (2021); Liu et al. (2023b), where agents share a joint reward according to their contributions, and inter-agent communication Liu et al. (2023a); Guo et al. (2023), where agents exchange information in order to coordinate their joint policy. The most challenging multi-agent environments are partially observable, typically by limiting agents’ information to their close proximity, and the achievement of their common goals is predicated upon coordination, which needs to involve some form of communication. In most research, the action-value function and inter-agent communication are issues addressed separately. Therefore, in this paper, we introduce communication into the very core of the action-value function decomposition.
The prevailing paradigm in cooperative MARL is Centralized Training and Decentralized Execution (CTDE), where the training algorithm utilizes the global state, but agents can only use their own local observations. We argue that in most challenging MARL tasks, this is insufficient. Moreover, in most challenging MARL tasks, some form of inter-agent communication is a necessity, which can be provided relatively easily (e.g., through WiFi). Therefore, we introduce a communication protocol that collects information from each agent in a central entity and broadcasts the combined information to all agents. Because information aggregation is based on attention, our proposed architecture requires a constant number of weights, plus an embedding, for each agent. It is easy to apply using readily available and efficient implementations of the Transformer encoder layer.
The contribution of this paper may be summarized as follows:
  • We introduce Multi-Agent Communication via Transformer on Agents’ States (MACTAS) module for inter-agent communication in MARL. This module allows the agent to share their state of consciousness.
  • We verify MACTAS experimentally on several challenging maps from SMACSamvelyan et al. (2019) and SMACv2 Ellis et al. (2023), giving excellent results and establishing state-of-the-art in some cases.
2. Related Work
Single-Agent Reinforcement Learning (SARL) Sutton and Barto (2018) addresses trial-and-error learning in sequential decision-making under uncertainty. In Multi-Agent Reinforcement Learning (MARL) Ning and Xie (2024), a group of agents learns to act in the same dynamic environment, collectively modifying its state. There are various settings for MARL, depending on whether the agents operate synchronously or asynchronously, whether they have common or adverse goals, and other circumstances. In this paper, we focus on the synchronous and cooperative MARL, with a single reward value for all agents at every time instance.
MARL has been addressed by extending well-established SARL algorithms, including Multi-Agent SAC Wei et al. (2018), Multi-Agent PPO Yu et al. (2022), and Multi-Agent DDPG Lowe et al. (2017).
Methods of cooperative MARL often adopt a centralized training with decentralized execution (CTDE) paradigm, in which agents share information during training but act based solely on their own observations during execution. However, this paradigm suffers from the global action-value function, whose complexity grows exponentially with the number of agents Kraemer and Banerjee (2016). Therefore, action-value function decomposition is a fundamental problem in MARL. Existing decomposed architectures include modules assigned to agents. These modules are fed with agents’ observations and their previous actions and produce agents’ contributions to the architecture’s output. The best action for the agent is the one that makes the highest contribution. In VDN Sunehag et al. (2018), these contributions are just added. In QMIX Rashid et al. (2018), a neural module, the mixing network, produces its weighted sum with positive weights. In Weighted QMIX Rashid et al. (2020), the architecture is trained with a special weight assigned to the collective actions deemed the best. In QTRAN Son et al. (2019), an additional module fed with all the actions of all agents corrects the weighted sum of the agents’ contributions. In Qatten Yang et al. (2020), the mixing network is based on attention. In CDS Li et al. (2021), a component of the mixing network optimizes the amount of exploration undertaken by the agents. In DVD Li et al. (2022a), the internal state of the modules assigned to agents is also passed to the mixing networks. In SHAQ Wang et al. (2022), the mixing network is enriched by several components based on Shapley value theory. In QPLEX Wang et al. (2021), a carefully engineered duplex dueling of pairs of contributions of agents is introduced into the mixing network. In NA2Q Liu et al. (2023b), a similar duplex dueling is performed with the use of attention.
Another fundamental problem in cooperative MARL is the coordination of agents’ activity for which a certain communication protocol is required. A number of solutions have been proposed for this purpose: CommNet Sukhbaatar et al. (2016) introduces differentiable communication by aggregating a learned broadcast vector from each agent, TarMAC Das et al. (2019) uses signature-based soft attention to direct agent communication towards recipients, NDQ Wang et al. (2019) applies two forms of regularization – mutual information maximization and message entropy minimization to optimize message content, GA-Comm Liu et al. (2020) uses two-stage attention mechanism to limit number of messages exchanged between agents, IS Kim et al. (2021) encodes trajectory of future actions for each agent and shares them with others, in MAIC Yuan et al. (2022) agents explicitly model their teammates and use it to bias their value function, DHCG Liu et al. (2023a) learns a dependency graph between agents and uses it to limit the number of communication edges, TEM Guo et al. (2023) uses a Transformer to simulate communication resembling email exchange, where agents only directly communicate with teammates nearby, but can forward messages from one to another, ACE Li et al. (2023) proposes bidirectional action-dependent Q-learning.
In addition, there are solutions in which the agents communicate in natural language, including: TWOSOME Tan et al. (2024) which uses a large language model (LLM) to evaluate the probabilities of joint action and applies LoRa for efficient PPO training, and Verco Li et al. (2024) which separates the output of the LLM into communication and actions, allowing LLM agents to exchange interpretable verbal messages.
We acknowledge various applications of attention or Transformer architectures in MARL algorithms, such as in Qatten mixing network, or in communication algorithms like NA2Q, GA-Comm, and TEM. None of the previous works has applied the Transformer architecture directly as a communication protocol. Our work fills this gap. We consider communication methods that can be coupled with an arbitrary action-value decomposition method. We also allow agents to send messages based only on their local information (current observation, historical observations, and historical actions). In this framework, the best results so far have been achieved by MAIC Yuan et al. (2022).
3. Method
Figure 2. The proposed MACTAS architecture.
3.1. Problem definition
In this paper, we discuss the problem of MARL using the formalism of a cooperative Decentralized Partially Observable Markov Decision Process (Dec-POMDP) with communication. Dec-POMDP is defined by a tuple $\langle\mathbb{N},\mathbb{S},\mathbb{A},\mathcal{R},\mathcal{P},\Omega,\mathcal{O},\gamma\rangle$, where $\mathbb{N}=\{1,\dots,n\}$ represents the team of agents, $\mathbb{S}$ is the global environmental state space, $\mathbb{A}=\mathbb{A}^{1}\times\dots\times\mathbb{A}^{n}$ is the joint action space, $\mathcal{R}$ is the reward function, $\mathcal{P}$ is the state transition probability function, $\Omega=\Omega^{1}\times\dots\times\Omega^{n}$ is the joint observation space, $\mathcal{O}$ is the observation function and $\gamma$ is the discount factor. At time $t=1,2,\dots$ each $i$-th agent makes an observation, $o^{i}_{t}=\mathcal{O}^{i}(s_{t}),o^{i}_{t}\in\Omega^{i}$, of the environment state $s_{t}\in\mathbb{S}$. Agents may exchange messages with others through a certain channel that may generally refine these messages with some processing.
Each agent maintains a representation, $\mathcal{I}^{i}_{t}$, of observations it has made, messages it has received, and actions it has taken. Based on thereof and a policy, $\pi^{i}$, the agent performs an action, $a^{i}_{t}\in\mathbb{A}^{i}$, $a^{i}_{t}\sim\pi^{i}(\cdot|\mathcal{I}^{i}_{t})$. The team action $a_{t}=[(a^{1}_{t}),\dots,(a^{n}_{t})]$ impacts the next environment state $s_{t+1}\sim\mathcal{P}(\cdot|s_{t},a_{t})$ and the reward $r_{t}=\mathcal{R}(s_{t},a_{t},s_{t+1})$ is given collectively to all agents.
The goal of the agents’ learning is to optimize their team policy $\pi=[\pi^{1},\dots,\pi^{n}]$ to expect in each environment state $s$ the highest sum of discounted rewards
$\mathcal{V}^{\pi}(s)=\mathbb{E}\left(\sum_{i\geq 0}\gamma^{i}r_{t+i}\Big|s_{t}=s,\text{policy in use}=\pi\right),$
where $\gamma\in[0,1]$.
The goal of this paper is to establish the message exchange protocol that supports the above goal of learning.
3.2. MACTAS
Our proposed Multi-Agent Communication via Transformer on Agents States (MACTAS) architecture is presented in Fig. 2. It is built on a decomposition of the action-value function. Each such decomposition includes the input modules assigned to the agents ($f^{i}$ assigned to the $i$-th agent for $i=1,\dots,n$ in Figure 2). These modules are fed with previous actions of these agents, $a^{i}_{t-1}$, and their current observations, $o^{i}_{t}$. They produce agents’ contributions, $\widehat{Q}^{i}_{t}$, to future discounted rewards and perhaps some other vectors Li et al. (2022a). The outputs of the $f^{i}$ modules are fed to the mixing network that produces the value of the action-value function, $\widehat{Q}_{t}$. The input module $f^{i}$ usually comprises a feedforward part, denoted by MLP in Figure 2, a recurrent part, denoted by GRU, and the output feedforward part, again denoted by MLP.
Intuitively, we understand the $f^{i}$ module as the consciousness of the $i$-th agent. Based on what the agent is aware of, it determines its contribution to future discounted rewards, depending on the action it currently takes. In order to give a global view on the current state to the agent’s consciousness, we introduce a communication mechanism. We divide the $f^{i}$ module after its recurrent part and feed its state $h^{i}_{t}\in\real^{n_{h}}$ to the module $g$, where the states of consciousness of the agents are confronted. As a result, the $g$ module produces increments, $z^{i}_{t}\in\real^{n_{h}}$. The sum $z^{i}_{t}+h^{i}_{t}$ is then passed to the second MLP part of the $f^{i}$ module.
Consequently, the consciousness of each agent is based not only on its observations and previous actions, but also on observations and previous actions of all other agents.
The minimal requirement for the $g$ module is as follows:
  1. (1)
    It is fed with $n$ vectors and produces $n$ vectors, both of size $\dim(h^{i}_{t})$. $i$-th input vector corresponds to the $i$-th output vector.
  2. (2)
    It is a layered structure and the layers include self-attention Vaswani et al. (2017).
Introducing $g$ that satisfies the above requirements in their basic form is enough to outperform MARL without communication by a large margin. However, what follows from our experiments is that usually the best performing form of the $g$ module is a stack of Transformer (encoder) modules Vaswani et al. (2017), with the following settings:
  1. (1)
    There are no positional embeddings. The agents’ embeddings are fed to $f^{i}$ instead.
  2. (2)
    There is no masking.
  3. (3)
    We use a separate Adam optimizer Kingma and Ba (2017) for the Transformer parameters; the rest of the architecture is trained with its originally proposed optimizer (typically it is RMSProp Tieleman and Hinton (2012)).
  4. (4)
    The weights of the output layer of the top Transformer module are initialized with zeros. Therefore, initially, this module does not impact the state of consciousness of the agents.
A good performance of the Transformer in this role is an expected result. This architecture has been specifically optimized to transform a set of embeddings into another set of the same size, which is exactly what we require in the $g$ module.
3.3. Implementation considerations
The basic approach to implement a multi-agent system with our proposed MACTAS module is to set up the Intermediate Processing Unit (IPU) on which the module is run. At each time instant, all $n$ agents transmit their internal states $h^{i}_{t}$ to the IPU. Then, the Transformer layers process all the input vectors and transmit the increments $z^{i}_{t}$ back to the agents. The number of operations necessary to process the Transformer modules is $O(n^{2})$. However, they are very efficiently parallelizable. The total size of all the transmissions is $O(n)$, specifically, equal to $2n\dim(h^{i}_{t})$.
Usually in multi-agent systems, reducing the communication capacity to $O(n)$ and offloading the communication processing burden to the IPU is very convenient. In these systems, the agents are naturally connected through a network and their computational capacity is limited (e.g., they are mobile and battery-powered), contrary to that of the network infrastructure. Still, these are the agents that need to take on the processing of their observations and control.
In most inter-agent communication protocols, including MAIC Yuan et al. (2022), all pairs of agents need to be able to communicate, which translates into $O(n^{2})$ messages. Technically, the parts of the agents that communicate with others could be moved to an IPU. Communication through the IPU would require $O(n)$ messages from the agents to the IPU and back. Still, natural parallelization of MACTAS is its noticeable advantage.
A distributed implementation of a multi-agent system with our proposed module is also possible. Then, each agent needs to take on the part of processing that consists of refining $h^{i}_{t}$ and results in $z^{i}_{t}$. That requires exchange of messages with other agents as many times as many Transformer layers are stacked in MACTAS, each time with $O(n)$ messages; $O(n^{2})$ for all agents. In training, the agents need to aggregate increments in their shared weights.
In the distributed implementation, it may happen that some of the agents are out of reach of others and do not exchange information with them. That only means that self-attention in MACTAS is forced to assign zero weight to unavailable transmissions from those unheard agents.
3.4. Exploration
Learning in a multi-agent system with decentralized execution is hampered by agents’ policies falling into suboptimal Nash equilibria. Let us consider a typical MARL setting where agents use the $\epsilon$-greedy strategy with $\epsilon$ annealing and explore independently. On average, only $n\epsilon$ agents explore. To learn that a certain combination of agents’ actions is better than the current Nash equilibrium, at least two agents need to explore and try a better combination. However, this happens rarely. What happens more often is that only one agent explores, which leads to a combination worse than greedy. Note that this effect may be even stronger for a better approximation of the $Q$ function, as it leaves less room for accidental switches of the greedy action combinations.
To counteract this effect, we propose a combination of $\epsilon$-greedy and Boltzmann exploration. Namely, each agent with probability $\epsilon$ chooses independently a random action; however, with probability $1-\epsilon$ it does not act greedily, but chooses from the top $k$ of its actions (according to its local action-value function) using the Boltzmann strategy. Using $k=1$ is equivalent to the pure $\epsilon$-greedy policy. Selecting a small $k$ leads to exploring a region that is quite good according to the agent. We observe that $k=2$ works well in practice. The temperature of the Boltzmann exploration controls the amount of extra exploration that we want to add.
Hyperparameter Value
Batch size 32
Test episodes 32
Replay buffer size 5000
Discount factor 0.99
Start epsilon 1.0
Finish epsilon 0.05
Anneal steps 50,0001 or 100,0002
RNN units 64
Optimizer RMSProp
Learning rate 0.0005
Table 1. Basic experimental settings and hyperparameters. 1 for 2.5 mln. step maps. 2 for 7.5 mln. step maps.
Scenario Stacked encoders FFN dimension
2c_vs_64zg 2 512
5m_vs_6m 1 128
27m_vs_30m 1 128
MMM2 3 256
6h_vs_8z 4 512
3s5z_vs_3s6z 3 512
All SMACv2 scenarios 3 512
Table 2. MACTAS shape for each scenario.
Figure 3. Percentage of wins in test games in training time for the connection protocols MACTAS and MAIC, and the $Q$ architectures QMIX, QPLEX, and VDN.
Figure 4. Percentage of wins and standard deviations in test games in training time for the connection protocols MACTAS, MAIC, and the bare mixer for the $Q$ architecture QMIX
Figure 5. The impact of our exploration scheme on MACTAS+QMIX.
Figure 6. The impact of our residual connection on MACTAS+QMIX.
4. Experiments
In this section, we present the results of our proposed MACTAS communication protocol in the StarCraft Multi-Agent Challenge environments (SMAC Samvelyan et al. (2019) and SMACv2 Ellis et al. (2023)). We compare it with the state-of-the-art MAIC Yuan et al. (2022), in combinations with 3 mixer algorithms: QMIX, QPLEX, and VDN, on 6 maps. We present results of two ablation studies regarding the design of neural network connectivity and the exploration form. Additionally, in the Appendix, we report the sensitivity study of MACTAS to hyperparameters, lower-level SMAC and SMACv2 settings, and analyses of the most interesting strategies that have emerged in training.
4.1. StarCraft Multi-Agent Challenge
The StarCraft Multi-Agent Challenge (SMAC and SMACv2) is a micromanagement benchmark where two opposing teams face each other in different scenarios (maps), with one team controlled by the RL agent and the other by a built-in algorithm. Maps differ by difficulty, mainly due to an imbalance in the number of controlled units. For our experiments, we selected 2 hard and 4 super-hard SMAC scenarios, which are widely benchmarked in the MARL community:
  • 2c_vs_64zg with 2 colossus units leveraging terrain advantage against 64 small enemy zerglings. Level: hard.
  • 5m_vs_6m with 5 cooperating marines and 6 enemy marines. Level: hard.
  • 27m_vs_30m big battle of 27 marines conquering 30 marines. Level: super-hard.
  • MMM2 scenario with diverse units containing 1 medivac (medic), 2 marauders, and 7 marines fighting team with 1 more marauder and 1 more marine. Level: super-hard.
  • 6h_vs_8z with 6 shooting hydralisks against 8 malee zealots. Level: super-hard.
  • 3s5z_vs_3s6z scenario with 3 stalkers and 5 zealots against 3 stalkers and 6 zealots. Level: super-hard.
To verify the versatility of the MACTAS algorithm, we run experiments on a diverse set of hard and super-hard scenarios. Every map provides the opportunity to find a unique winning strategy. The agent requires learning basic skills, such as navigating the battlefield, managing health and shield, or focusing fire. All considered scenarios are imbalanced in favor of the enemies; therefore, simple strategies are insufficient to win the episode. Popular observed winning policies include crossing the obstacle on 2c_vs_64zg, flanking on 27m_vs_30m, ’kiting’ on 6h_vs_8z, ’hit and run’ on 3s5z_vs_3s6z Samvelyan et al. (2019).
In addition to classical SMAC, we benchmark the algorithms in its newer iteration, SMACv2 Ellis et al. (2023), which introduces new maps focused on scenario randomization. Each of the new maps defines the number of units in each team and the race (the same for both teams), while the exact types and starting positions are randomized according to the configuration. We use the default configuration for each scenario recommended in the SMACv2 repository. Our experiments focus on more challenging configurations, where the player team is outnumbered by the enemy team controlled by in-game AI, with 6 tested scenarios: zerg_5_vs_5, zerg_10_vs_11, zerg_20_vs_23, protoss_5_vs_5, protoss_10_vs_11, and protoss_20_vs_23.
4.2. Experimental setting
We borrow the basic settings from NA2Q Liu et al. (2023b) and extend their code with the implementation of MACTAS and MAIC. The algorithms are trained using the Double Q-Learning Hasselt et al. (2016) algorithm. Each experiment is run for a fixed number of training episode steps – 2.5 million for 2c_vs_64zg, 5m_vs_6m, MMM2, 27m_vs_30m environments and 7.5 million for 6h_vs_8z, 3s5z_vs_3s6z and all SMACv2 environments, interleaved with test episodes every 2000 steps. Each experiment is repeated for seeds $1,\dots,5$. We report the average percentage of wins in test games, along with the standard deviations in the Appendix.
All hyperparameters not strictly related to the communication algorithm are shared between maps, algorithms, and mixers and are presented in Table 1.
For MAIC communication, we employ the same hyperparameters as those used in the original paper.
MACTAS hyperparameters
We notice that MACTAS is sensitive to communication hyperparameters depending on the scenario setup. The parameters are the number of stacked Transformer encoders and the dimension of their positional feedforward network block. We theorize that the depth and width of the communication channel create a trade-off between the ability to learn a complex joint policy (eg. kiting, hit-and-run), and the time it takes to learn an effective policy.
We perform a hyperparameter search for each scenario, for which detailed results are included in the Appendix, and use the values presented in Table 2.
Mixer compatibility
Both MACTAS and MAIC are agnostic to the value function decomposition algorithm, so we compare both algorithms in combination with commonly used mixers: VDN, QMIX, and QPLEX. Each combination named MACTAS+VDN, MAIC+QPLEX, etc., is tested across all maps and compared separately to provide a fair comparison and isolate the effect of the mixer algorithm.
4.3. Results
SMAC
The basic results are presented in Fig. 3. We compare MACTAS with MAIC on 6 maps and 3 mixers, which makes 18 combinations. If we consider the area under the learning curve, MACTAS performs better than MAIC in 9 cases, MAIC performs better in 3 cases and there are 6 draws. If we consider the final performance, MACTAS yields better results 6 times, MAIC yields better results 3 times, and there are 9 draws.
On the 2c_vs_64zg map, all algorithms except MAIC+VDN achieve the same final performance. In terms of the area under the learning curve MACTAS performs similarly to MAIC for the QMIX and VDN mixers and better for the QPLEX mixer. MACTAS+QMIX and MAIC+QMIX give jointly the best results on this map. On the 5m_vs_6m map MACTAS+QPLEX achieves the same final result as MAIC+QPLEX but with a better area under the learning curve. MACTAS+QMIX gives the same final result as MAIC+QMIX, but has a worse area under the learning curve (MAIC+QMIX drops around two million steps). MACTAS+VDN achieves a better final result than MAIC+VDN with a similar area under the learning curve (MAIC+VDN starts to learn faster but is unable to go beyond $60\%$ of the mean of the won battles). On this map MACTAS+QPLEX gives the best result. We observe that on the 27m_vs_30m map, our algorithm achieves a better area under the learning curve for all three mixers, and for QPLEX, it also achieves a better final performance. For the other mixers, the final performance is similar. We hypothesize that the good performance of our algorithm is because the number of parameters in our algorithm is independent of the number of agents. Thus, our method scales well in environments with a large number of agents. Overall, MACTAS+QMIX gives the best result on this map. On the MMM2 map, MAIC performs better for all three mixers in terms of the final result. However, the areas under the learning curves are similar and for MACTAS+QMIX and for MAIC+QMIX, as our algorithm starts to learn faster. MAIC+QPLEX performs the best on this map. This is the only map on which a version of our algorithm does not give the best result. On the 6h_vs_8z map MACTAS performs better than MAIC for the QMIX mixer in terms of the area under the learning curve and the final result. MACTAS+VDN gives a similar final result as MAIC+VDN, but with a better area under the learning curve. MACTAS+QPLEX achieves a similar final result as MAIC+QPLEX, but with a worse area under the learning curve. Overall, MACTAS+QMIX and MAIC+QPLEX give jointly the best result on this map. On the 3s5z_vs_3s6z map, MACTAS+QMIX and MACTAS+VDN give better results in terms of both the area under the learning curve and the final result than MAIC+QMIX and MAIC+VDN, respectively. MACTAS+QPLEX gives a similar final result to MAIC+QPLEX, but with a worse area under the learning curve. MACTAS+VDN performs the best on this map.
SMACv2
The basic results are presented in Fig. 4. We compare MACTAS with MAIC and the bare mixer on 6 maps and the QMIX mixer. In all scenarios, our algorithm outperforms MAIC and the bare mixer. We observe that the differences between MACTAS+QMIX and MAIC+QMIX are similar to the differences between MAIC+QMIX and the bare mixer.
4.4. Ablations
There are two objectives of the ablation study. Firstly, we notice that the Transformer module has residual connections inside it. The question is whether the residual connection around it (the arrow between GRU and “++” in Fig. 2) is helpful indeed. Figure 6 shows the impact of the residual connection on MACTAS+QMIX. We observe that the residual connection is crucial for the 3s5z_vs_3s6z and 6h_vs_8z maps.
Secondly, we verify the role of our proposed exploration scheme. Figure 5 shows the impact of our exploration scheme for MACTAS+QMIX. We observe that our exploration method improves the results on the most challenging SMAC maps, which require $7.5$ million training steps: 3s5z_vs_3s6z, and 6h_8z. It is also indifferent on the maps: 2c_vs_64zg, 27m_vs_30m, and MMM2. Our exploration scheme slows down the training for the 5m_vs_6m map. We used softmax temperature $0.33$ for the map 3s5z_vs_3s6z and $0.25$ for the maps: 2c_vs_64zg, 5m_vs_6m, 27m_vs_30m, MMM2, and 6h_8z.
4.5. Discussion
[…]
The modular nature of MACTAS is its great advantage. There are widespread variants of the Transformer code optimized for different machines. This makes the implementation and deployment of MACTAS in a given multi-agent system straightforward.
Limitations
A moderate sensitivity of the performance of our proposed communication module to its shape requires searching for its best setting in additional runs. This slightly slows down the deployment of MACTAS.
5. Conclusions and Future Work
[…]
We observe that $\epsilon$-greedy exploration applied independently by agents in MARL leaves room for improvement since it condones the agents’ policies falling into suboptimal Nash equilibria. We addressed this issue by proposing a combination of $\epsilon$-greedy and Boltzmann exploration, which increased the performance of MACTAS+QMIX on the 6h_vs_8z, and the 3s5z_vs_3s6z SMAC maps. However, we believe that this issue requires further research and intend to undertake it in the future.
{acks}
We gratefully acknowledge the Polish high-performance computing infrastructure PLGrid (HPC Center: ACK Cyfronet AGH) for providing computer facilities and support within the computational grant no. PLG/2025/018560 and no. PLG/2025/017992. We thank the University of Warsaw for providing access to the computing infrastructure.
Appendix A StarCraft Multi-Agent Challenge (SMAC & SMACv2)
As a benchmark for MACTAS, we chose the StarCraft Multi-Agent Challenge (SMAC) Samvelyan et al. (2019) and its newer version SMACv2 Ellis et al. (2023), which are simplified versions of the StarCraft II game. The framework exists in two versions – SMAC (code available at https://github.com/oxwhirl/smac), which introduced predefined deterministic scenarios that often include purposeful selection of units or obstacles placement, and SMACv2 (code available at https://github.com/oxwhirl/smacv2), which added new randomized scenarios based on plain maps and randomized units generation. From here onward, we will use ”SMACv1” to refer to the first version of the benchmark and ”SMAC” to refer to both SMACv1 and SMACv2, making an explicit distinction where necessary. SMAC uses PySc2 to communicate with the game engine https://github.com/google-deepmind/pysc2. SMAC is characterized by the decentralized management of each unit, where each unit acts as an independent agent. This feature allows for the testing of decentralized algorithms, including those originating from MARL. Benchmark scenarios are limited to combat with a limited number of units. Those scenarios do not account for the economic aspect of the game, the so-called “macro” part of the game – collecting resources and building structures. Resources in the game (gas and minerals) are required to build units and structures. Structures enable the recruitment of units and their subsequent upgrading. The considered scenarios do not involve “worker” units, which collect resources. Instead, the scenarios focus on ’micro’, a term used in the StarCraft II community to describe unit control during combat. In SMACv2, the new addition is randomized maps defined by race (same for both teams), number of units in each team, thus map names like zerg_10_vs_11. The new maps randomize types of units according to the probability of each unit type defined in scenario configuration files called capabilities config. In addition, the configuration allows for specifying the units’ start position, their spread, and configuration – either surround, reflect, or random probability of each. The challenge for a human player is to multitask between these two tasks – developing the game economy and controlling the fighting units. SMAC focuses solely on combat, thereby ensuring repeatability of this benchmark, as each iteration involves a fight between the same specification units, eliminating uncertainty due to unit upgrades and different strategies in the economic part of the game. In StarCraft II, there are three races, and our experiments covered controlling units of all of them. The process for running SMAC is described in our repository.
Appendix B Detailed Results
Figures 7, 8, and 9 contain more detailed results of our experiments described in Section 4.3. The curves on the plots represent the mean percentage test game win rate during the algorithm training for 6 different environments and various mixers. The shaded area around the curves is the standard deviation of the results among 5 different seeds (1 to 5) of the random number generator. For each seed, the mean percentage test win rate was calculated based on 32 independent test episodes. We observe robustness of MACTAS with the QMIX mixer – on the most challenging environments 6h_vs_8z and 3s5z_vs_3s6z, it outperforms MAIC+QMIX by a large margin; on 27m_vs_30m, it learns much faster; on 2c_vs_64zg and 3s5z_vs_3s6z, it performs similarly, and only slightly worse on the MMM2 map. For the QPLEX mixer, our method performs much better than MAIC on the 2c_vs_64zg and 27m_vs_30m maps, slightly better on the 5m_vs_6m, slightly worse on MMM2, learns slower but achieves similar final result on the 6h_vs_8z and 3s5z_vs_3s6z maps. Our method combined with the VDN mixer gives significantly better results than MAIC on the 3s5z_vs_3s6z map, slightly better on 6h_vs_8z, 27m_vs_30m, 5m_vs_6m, marginally better on the 2c_vs_64zg map, and marginally worse on the MMM2 map.
Figure 7. Means and standard deviations of the percentage wins in the test games during training for QMIX with MAIC, MACTAS, and bare.
Figure 8. Means and standard deviations of the percentage wins in the test games during training for QPLEX with MAIC, MACTAS, and bare.
Figure 9. Means and standard deviations of the percentage wins in the test games during training for VDN with MAIC, MACTAS, and bare.
Appendix C Tactics used by MACTAS
The SMAC challenge forces agents to learn various tactics, such as attacking from a distance with melee units (’kiting’) or using a movement advantage (’hit and run’). In this section, we provide an overview of the tactics used by our model based on our understanding of StarCraft II. We observed differences in the behavior of our model depending on the random seed. The description of tactics is based on replays of our models for a seed value of 1.
C.1. 2c_vs_64zg
Figure 10. 2c_vs_64zg replay.
In this scenario, agents control 2 colossi units. Colossi can move through hills between the high and the low ground. Against them, 64 enemy zerglings are deployed, which are fast ground units. Zerglings are most effective when they surround an opposite unit. In the scenario 2c_vs_64zg, the colossi start on the hill, while the zerglings are split into two groups of 32 on the left and the right, both of them on the low ground. Figure 10 presents the following tactics: colossi attack the right group of zerglings and wait for the left group to climb the hill. Then, colossi use the ability to traverse the hill and attack the remaining zerglings, which are forced to run around the hill. In this way, the colossi avoid being surrounded and thereby defeated.
C.2. 27m_vs_30m and 5m_vs_6m
Figure 11. 27m_vs_30m replay.
Figure 12. 5m_vs_6m replay.
The 27m_vs_30m and the 5m_vs_6m scenarios consist of marine units only. They are similar in the way that the opposing force numerically outnumbers the agents under control. In Figure 11, the agent’s units place themselves in a semicircle formation, also known as flanking, which maximizes the number of shooting units, while the enemies block each other. Units controlled by MACTAS focus fire on enemy units with lower health to quickly reduce the number of opponents. Furthermore, units with low health retreat and continue fighting, while avoiding being shot. This behavior is presented in Figure 12: Agents with low health retreat when others focus fire on one enemy unit. The winning strategy is similar for both 5m_vs_6m and 27m_vs_30m. 5m_vs_6m is labeled as a hard benchmark, while 27m_vs_30m is classified as super-hard. The percentage imbalance between units is higher in 5m_vs_6m. However, 27m_vs_30m has a larger number of units that need to cooperate.
C.3. MMM2
Figure 13. MMM2 replay.
The MMM2 is a scenario with a diverse set of units containing 1 medivac (medic), 2 marauders, and 7 marines fighting against a team with 1 more marauder and 1 more marine. Medivac cannot attack, but can heal units as long as it has energy. MACTAS ground units focus fire on enemy units and eliminate them one by one, and fall back when their health is low. The enemy medivac does not change healed unit, while the MACTAS’ medivac changes healed unit to the one with a critically low health or to the currently attacked unit. Units trained with MACTAS move medivac to the front line and focus enemy fire on it, and when its health level is low, retreat it. MACTAS’ medivac taking enemy fire secures outnumbered marines and marauders. Typically, human players try to save their medivacs because they are used for healing units between fights during a regular game. Only under certain conditions do they sacrifice their medivacs; for example, when they are breaking through an opponent’s line of defense to focus enemy fire on the medivacs and away from damage-dealing units. In this scenario, we can observe similar behavior in the MACTAS medivac control, which results in winning combat against an opponent who has a numerical advantage.
C.4. 6h_vs_8z
Figure 14. 6h_vs_8z replay.
Agents control 6 range shooting units, hydralisks, while the enemy controls 8 melee zealots. In Figure 14, we can see a tactic called “kiting” – hydralisks split into two groups and are attacking zealots from distances. Enemy zealots attack the nearest unit – our agents learned to move enough to change the focus of approaching units and then continue striking. Notably, the zealot units within this scenario are not upgraded with the “charge” ability, which would otherwise allow them to have a significant increase in movement speed and easily reach hydralisks.
C.5. 3s5z_vs_3s6z
Figure 15. 3s5z_vs_3s6z replay.
MACTAS controls 3 stalkers and 5 zealots, while the enemy has one more zealot. The stalkers are fast, range-attacking units, but with weak defense. On the other hand, zealots have a good defense, but they need to get close to the opponent’s unit. The tactic in this scenario is interesting. Using one stalker as a decoy, MACTAS first defeats the enemy stalkers and then proceeds to eliminate the rest of the enemy force. One agent, a stalker, sacrifices himself and focuses a few enemy zealots on himself. In Figure 15, we can see 5 enemy zealots following the stalker marked as “Local Player”. The remaining MACTAS’ 2 stalkers and 5 zealots attack 3 enemy stalkers and 1 zealot with an advantage. This move results in the elimination of enemy stalkers and the zealot. In the end, MACTAS’ units eliminate zealots lured by the sacrificed stalker.
Appendix D Sensitivity analysis
Figure 16. MACTAS; sensitivity of the results to architecture size with the QMIX mixer.
Figure 17. MACTAS; sensitivity of the results to architecture size with the QPLEX mixer.
Figure 18. MACTAS; sensitivity of the results to architecture size with the VDN mixer.
Figure 19. MACTAS; sensitivity of the results to the dropout parameter in the Transformer block. The default value of this parameter is $10\%$.
Figure 20. MACTAS; sensitivity of the top 2 exploration methods to the temperature coefficient.
We evaluated the choice of the number of Transformer encoder layers and the width of the Transformer feedforward layer for MACTAS with QMIX, QPLEX, and VDN as a mixer on six SMAC(v1) maps: 2c_vs_64zg, 5m_vs_6m, 27m_vs_30m, MMM2, 6h_vs_8z, and 3s5z_vs_3s6z. We experimented with the dropout of the Transformer encoder on all six maps. We also assessed the sensitivity of our new top $k$ exploration technique to the softmax temperature in the case of MACTAS+QMIX and $k=2$ on all six maps.
[…]
Figure 19 shows how the dropout parameter affects the performance of MACTAS. We observe that the absence of dropout harms the performance; however, the choice of the dropout fraction is not crucial when set in a reasonable interval; relevant differences can be observed only on the 3s5z_vs_3s6z map.
Figure 20 shows the impact of the softmax temperature parameter on the proposed exploration scheme. We observe that this method is sensitive to this parameter. For the 2c_vs_64zg, 5m_vs_6m, 27m_vs_30m, MMM2, and 6h_vs_8z maps, the low temperatures ($0.25$, $0.33$) work best with better results for the $0.25$ temperature. We can observe an interesting situation on the 3s5z_vs_3s6z map. The best results are obtained for the $0.33$ temperature, with the temperature $4$ being the second-best, and the temperature $0.25$ performing quite poorly, which suggests that, in general, the relationship between the temperature value and performance may not be monotonic. For the most challenging maps, which require 7.5 million training steps, the top 2 exploration schemes outperform the $\epsilon$-greedy exploration, which is equivalent to choosing 0 as the temperature. In other cases, the new exploration scheme does not outperform the $\epsilon$-greedy exploration; however, except for the case of the 5m_vs_6m map, where training is slowed down, the method remains indifferent.
Appendix E Software
Table 3 contains a list of key programming libraries used during our experiments. The complete list of programming libraries is available in the MACTAS’ official code repository in the file requirements.txt.
Library name Version
Python 3.9.21 or 3.9.22
numpy 1.23.1
nvidia-cublas-cu12 12.4.5.8
nvidia-cuda-cupti-cu12 12.4.127
nvidia-cuda-nvrtc-cu12 12.4.127
nvidia-cuda-runtime-cu12 12.4.127
nvidia-cudnn-cu12 9.1.0.70
nvidia-cufft-cu12 11.2.1.3
nvidia-curand-cu12 10.3.5.147
nvidia-cusolver-cu12 11.6.1.9
nvidia-cusparse-cu12 12.3.1.170
nvidia-nccl-cu12 2.21.5
nvidia-nvjitlink-cu12 12.4.127
nvidia-nvtx-cu12 12.4.127
scipy 1.13.1
torch 2.5.1
Table 3. Essential libraries in programming environment.
Appendix F Hardware
Cluster name GPU device GPU driver version CPU device Operating system Total RAM
Cluster 1 NVIDIA A100 SXM4 80GB 535.154.05 AMD EPYC 7742 64-Core Processor Linux-5.15.0-1059-nvidia-x86_64-with-glibc2.35 1024 GB
Cluster 2 NVIDIA A100 SXM4 40GB 570.133.20 AMD EPYC 7742 64-Core Processor Linux-5.14.0-503.26.1.el9_5.x86_64-x86_64-with-glibc2.34 1024 GB
Cluster 3 NVIDIA A100 PCIe 80GB 565.57.01 AMD EPYC 7713 64-Core Processor Linux-6.8.0-64-generic-x86_64-with-glibc2.39 2048 GB
Table 4. Specification of the computational clusters used in the experiments.
Here we present the hardware used in our experiments. We utilized the NVIDIA Ampere GPU architecture across three computational clusters. The specific setup is presented in Table 4.